# TeleMem: Building Long-Term and Multimodal Memory for Agentic AI

### v2

@@ 2.2.2 Memory Reading
Input: Inital query $q$, $\texttt{LLM}(\cdot)$, $\mathcal{M}=\mathcal{M}_{e}\cup\mathcal{M}_{\mathrm{obj}}$
@@
5while not exceed max_iterations do
6   action, args = $\texttt{LLM}(H)$;
7
8 if action is video.retriever then
9
10    $q\leftarrow$ args ;
[…]
@@ 10    $q\leftarrow$ args ;
13 else if action is video.rag then
14    $q\leftarrow$ args ;
[…]
@@ 14    $q\leftarrow$ args ;
17 else if action is video.qa then
18    $q,t_{start},t_{end}\leftarrow\mathtt{args}$ ;
19      results $\leftarrow$ $\texttt{video.qa}(q,t_{start},t_{end})$;
@@ 18    $q,t_{start},t_{end}\leftarrow\mathtt{args}$ ;
21 else if action is finish then
@@
25 $H=H+[(\text{action},\text{args},\text{results})]$;
@@
28return $\texttt{LLM}(H)$.
Algorithm 1 Multimodal Memory Reading.

### v4

@@ 3.3.2 Reasoning Paradigm: ReAct-Style Multimodal Agent
Input: Initial query $q$, language model $\texttt{LLM}(\cdot)$, multimodal memory $\mathcal{M}=\mathcal{M}_{e}\cup\mathcal{M}_{\mathrm{obj}}$
@@
5while not exceeding max_iterations do
6
7 $(\text{action},\text{args})\leftarrow\texttt{LLM}(H)$;
8
9 if action = video.retrieval then
10    $q\leftarrow\text{args}$;
11    $\text{results}\leftarrow\texttt{video.retrieval}(q,\mathcal{M})$;
@@ 10    $q\leftarrow\text{args}$;
13 else if action = video.rag then
14    $q\leftarrow\text{args}$;
15    $\text{results}\leftarrow\texttt{video.rag}(q,\mathcal{M})$;
@@ 14    $q\leftarrow\text{args}$;
17 else if action = video.qa then
18    $(q,t_{\text{start}},t_{\text{end}})\leftarrow\text{args}$;
19    $\text{results}\leftarrow\texttt{video.qa}(q,t_{\text{start}},t_{\text{end}})$;
@@ 18    $(q,t_{\text{start}},t_{\text{end}})\leftarrow\text{args}$;
21 else if action = finish then
@@
25 $H\leftarrow H\;\|\;[(\text{action},\text{args},\text{results})]$;
@@
28return $\texttt{LLM}(H)$;
Algorithm 1 ReAct-Style Multimodal Memory Reading
@@ Acknowledgments.
Appendix A Example Memory Entries
This section provides representative JSON snippets illustrating how different types of memory nodes are serialized in our system. These examples are intended to clarify the unified schema, field semantics, and provenance information used by the memory graph, rather than to enumerate all possible attributes.
Figure 3: Example JSON snippet of a text profile memory entry.
Text profile memory.
Figure 3 shows an example of a text profile memory entry, which captures relatively stable attributes of a user or an agent. Such entries typically include a high-level semantic summary, a persistent identifier, and lightweight metadata (e.g., timestamps or session indices) that enable long-term tracking and consolidation across interactions.
Figure 4: Example JSON snippet of a text event memory entry.
Text event memory.
Figure 4 illustrates a text event memory entry corresponding to a single interaction or dialogue turn. In addition to the summarized semantic content, the entry records the original messages, roles, and temporal information, enabling traceability back to raw conversational context and supporting dependency-aware retrieval.
Figure 5: Example JSON snippet of a video event memory entry.
Video event memory.
Figure 5 presents a video event memory entry that encodes localized multimodal observations. Besides the textual caption, such entries may include aligned timestamps, clip references, or auxiliary annotations, which allow the system to ground high-level semantics in concrete visual evidence.
Figure 6: Example JSON snippet of an object memory entry.
Object memory.
Figure 6 shows an object memory entry representing a persistent entity tracked across video frames or scenes. Typical fields include object identity, appearance attributes, behavioral cues, and the first observed timestamp, supporting long-term entity consistency and temporal reasoning.